OpenStudy (ilovecomputers1):

The function f(x) = 10(5)^x represents the growth of a lizard population every year in a remote desert. Crista wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Crista's purposes?

OpenStudy (ilovecomputers1):

f(x) = 10(5)the x over 2 power

OpenStudy (ilovecomputers1):

f(x) = ten halves(5)x f(x) = 10(5)x f(x) = 10(5 to the one half power)2x